well i hav a question about 2x2+15x-8=0 the 2 next to x is an exponent sorry i dont know how to type exponents. my question is that i know i need to use the quadratic formula and the answer is x=-8 and x=0.5 but i dont know why i need to use the quadratic formula and why the answer is x=-8 and x=0.5 . thanks for your time please answer me back asap
Answer by checkley71(8403)

2x^2+15x-8=0
(2x-1)(x+8)=0
2x-1=0
2x=1
x=1/2 answer.
x+8=0
x=-8 answer.